OpenStudy (moonlitfate):

Combine 4/i - 6/(8-i)

OpenStudy (moonlitfate):

Problem looks like this : \[\frac{ 4 }{ i } - \frac{ 6 }{ 8-i }\] I know that (8+i) is the conjugate of (8-i) and that the top and bottom of the second fraction has to multiplied by that. I'm not sure what to do with the first fraction, though. Help would be appreciated.

OpenStudy (anonymous):

multiply by i to both numerator and denominator

OpenStudy (anonymous):

1/i = -i

OpenStudy (anonymous):

does it make sense?

OpenStudy (moonlitfate):

None of my answer have 8 as a denominator, so that isn't correct. :(

OpenStudy (moonlitfate):

*the answers

OpenStudy (anonymous):

\[[\frac{ 4 }{ i }*\frac{ i }{ i }]-[\frac{ 6 }{ 8-i }*\frac{ 8+i }{ 8+i } ]\] you are done

OpenStudy (anonymous):

@MoonlitFate can you do the rest?

OpenStudy (moonlitfate):

I thought that both denominators had to be equal to each other to combine them. Similar to just normal fractions. All of the available answers have equal denominators...

OpenStudy (anonymous):

no , you have to rationalise first

OpenStudy (anonymous):

even if you'll do that , things will become more complex

OpenStudy (moonlitfate):

Hm. So to answer the first part of the equation the answer would be: \[\frac{ 4i }{ i ^{2} } - \frac{ 48 + 6i }{ 65}\]

OpenStudy (anonymous):

i guess it is correct

OpenStudy (anonymous):

\[i^2= -1\]

OpenStudy (anonymous):

\[-4i +\frac{ 48+6i }{ 65 }\] now it's just simple math

OpenStudy (moonlitfate):

Oh. Multiply the first term by 65 and just add. Got it. So the answer would be: \[\frac{ -48 }{ 65 } - \frac{ -266i }{ 65 }\]

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

i am sorry...you are absolutely right :)

OpenStudy (moonlitfate):

Ah, haha. I guess what confused me was one of the first steps. Why you had to multiply the first term by i^2